# Powerformer: A Section-adaptive Transformer for Power Flow Adjustment

### v4

@@ Abstract.
In this paper, we present a novel transformer architecture tailored for learning robust power system state representations, which strives to optimize power dispatch for the power flow adjustment across different transmission sections. Specifically, our proposed approach, named Powerformer, develops a dedicated section-adaptive attention mechanism, separating itself from the self-attention employed in conventional transformers. This mechanism effectively integrates power system states with transmission section information, which facilitates the development of robust state representations. Furthermore, by considering the graph topology of power system and the electrical attributes of bus nodes, we introduce two customized strategies to further enhance the expressiveness: graph neural network propagation and multi-factor attention mechanism. Extensive evaluations are conducted on three power system scenarios, including the IEEE 118-bus system, a realistic China 300-bus system, and a large-scale European system with 9241 buses, where Powerformer demonstrates its superior performance over several popular baseline methods.
@@ 1. Introduction
Recently, deep reinforcement learning (DRL) methods including DQN (Mnih et al., 2015), PPO (Schulman et al., 2017), and A2C (Mnih et al., 2016) have gained attention for their potential in addressing various control challenges within power systems, such as voltage control (Mu et al., 2024; Yang et al., 2020; Cao et al., 2022), economic dispatch (Hu et al., 2024; Dai et al., 2020; Li et al., 2021), and power flow adjustment (Wu et al., 2024; Gao et al., 2021; Wang et al., 2022). For the specific issue of power flow adjustment, these data-driven approaches effectively utilizes the robust capabilities of the pertained deep neural networks, enabling the direct learning of response actions in an end-to-end fashion, effectively tackling the high-dimensional control problems and providing a powerful solution for power flow adjustment. However, among the existing DRL-based methods (Wu et al., 2024; Gao et al., 2021; Wang et al., 2022), there are two evident limitations in the state representations for transmission section power flow adjustment:
@@
In this paper, we strive to develop an advanced transformer architecture for learning robust power system state representations, enabling the DRL agent to perform precise power dispatch strategies that stabilize the power system under various section settings. To tackle the first limitation, we develop a section-adaptive attention mechanism to enable an efficient amalgamation of section information with the power system state, leveraging the comprehensive capabilities of the transformer architecture (Vaswani et al., 2017; Zhu et al., 2023; Zhang et al., 2022; Ding et al., 2024) for the specific sectional system analysis. For the second limitation, we begin by analyzing the graph topology of the power system, and employing graph neural networks (Jing, 2023; Li et al., 2023; Zhang et al., 2023) to propagate features for embedding structural information into state representations. Furthermore, we introduce a multi-factor attention mechanism to efficiently fuse each state factor, ensuring comprehensive and accurate extraction of system states. Finally, we utilize the proposed Powerformer, integrating the Dueling DQN (Wang et al., 2016) framework to achieve precise and efficient control of power flow across transmission sections, thereby leading to improvements in the overall system performance.
@@ 2.2. Sectional Power System Network
Figure 2. The illustration of the Transformer and Powerformer architectures. (a) Transformer architecture with the multi-head self-attention mechanism (MHA). (b) Powerformer architecture with the multi-factor section-adaptive attention mechanism (MFSA).
@@ 3.2. Powerformer for Sectional Power System
where $\bm{V}_{t}^{(l)}$ and $\widetilde{\bm{A}^{(l)}}$ are obtained from Eq. 5 and 9, $\widetilde{\bm{A}^{(l)}}_{t}\in\mathbb{R}^{n\times 1}$ the $t$-th column associated with $t$-th electrical factor. Symbol $\odot$ denotes the Hadamard product with broadcast mechanism. It is important to note that the left matrix side of the $\odot$ operation is in $\mathbb{R}^{d\times n}$, while the right matrix is in $\mathbb{R}^{n\times 1}$. The resulting matrix after the Hadamard product will also reside in the space $\mathbb{R}^{d\times n}$.
@@ 3.3. The Implementation of Powerformer
Joint Training. To deliver the final dispatch policy based on the proposed Powerformer framework, we adopt the Dueling Deep Q Network (DQN) method [^1] The reason for selecting Dueling DQN is provide in Appendix E to train a deep reinforcement learning agent. Specifically, we introduce a Q-network $Q_{\theta}$ parameterized by $\theta$, which takes the state representation $\text{MFSA}(\mathcal{G})$ as input and estimates the Q-values for each possible dispatch action $d$. The dispatch policy with respect to $Q_{\theta}$ is defined as $\pi_{Q}(\mathcal{G})=\arg\max_{d\in\mathcal{D}}Q_{\theta}(\text{MFSA}(\mathcal{G}),d)$, where $\mathcal{D}$ is the dispatch action space.
@@ Appendix E The reason for selecting Dueling DQN
Appendix F More Ablation Studies on Three Datasets

### v5

@@ Abstract.
In this paper, we present a novel transformer architecture tailored for learning robust power system state representations, which strives to optimize power dispatch for the power flow adjustment across different transmission sections. Specifically, our proposed approach, named Powerformer, develops a dedicated section-adaptive attention mechanism, separating itself from the self-attention employed in conventional transformers. This mechanism effectively integrates power system states with transmission section information, which facilitates the development of robust state representations. Furthermore, by considering the graph topology of power system and the electrical attributes of bus nodes, we introduce two customized strategies to further enhance the expressiveness: graph neural network propagation and multi-factor attention mechanism. Extensive evaluations are conducted on three power system scenarios, including the IEEE 118-bus system, a realistic China 300-bus system, and a large-scale European system with 9241 buses, where Powerformer demonstrates its superior performance over several popular baseline methods. The code is available at: https://github.com/Cra2yDavid/Powerformer
@@ 1. Introduction
Recently, with the advancement of deep learning (LeCun et al., 2015; Shi et al., 2024; Vaswani et al., 2017; Shi et al., 2022, 2021), deep reinforcement learning (DRL) methods including DQN (Mnih et al., 2015), PPO (Schulman et al., 2017), and A2C (Mnih et al., 2016) have gained attention for their potential in addressing various control challenges within power systems, such as voltage control (Mu et al., 2024; Yang et al., 2020; Cao et al., 2022), economic dispatch (Hu et al., 2024; Dai et al., 2020; Li et al., 2021), and power flow adjustment (Wu et al., 2024; Gao et al., 2021; Wang et al., 2022). For the specific issue of power flow adjustment, these data-driven approaches effectively utilizes the robust capabilities of the pertained deep neural networks, enabling the direct learning of response actions in an end-to-end fashion, effectively tackling the high-dimensional control problems and providing a powerful solution for power flow adjustment. However, among the existing DRL-based methods (Wu et al., 2024; Gao et al., 2021; Wang et al., 2022; Wei et al., 2023; Qing et al., 2024), there are two evident limitations in the state representations for transmission section power flow adjustment:
@@
In this paper, we strive to develop an advanced transformer architecture for learning robust power system state representations, enabling the DRL agent to perform precise power dispatch strategies that stabilize the power system under various section settings. To tackle the first limitation, we develop a section-adaptive attention mechanism to enable an efficient amalgamation of section information with the power system state, leveraging the comprehensive capabilities of the transformer architecture (Vaswani et al., 2017; Zhu et al., 2023; Zhang et al., 2022; Ding et al., 2024) for the specific sectional system analysis. For the second limitation, we begin by analyzing the graph topology of the power system, and employing graph neural networks (Jing, 2023; Li et al., 2023; Yu et al., 2024; Chen et al., 2023a) to propagate features for embedding structural information into state representations. Furthermore, we introduce a multi-factor attention mechanism to efficiently fuse each state factor, ensuring comprehensive and accurate extraction of system states. Finally, we utilize the proposed Powerformer, integrating the Dueling DQN (Wang et al., 2016) framework to achieve precise and efficient control of power flow across transmission sections, thereby leading to improvements in the overall system performance.
@@ 2.2. Sectional Power System Network
Figure 2. The illustration of the Transformer and Powerformer architectures. (a) Transformer architecture with the multi-head self-attention mechanism (MHA). (b) Powerformer architecture with the multi-factor section-adaptive attention (MFSA).
@@ 3.2. Powerformer for Sectional Power System
where $\bm{V}_{t}^{(l)}$ and $\widetilde{\bm{A}^{(l)}}$ are obtained from Eq. 5 and 9, $\widetilde{\bm{A}^{(l)}}_{t}\in\mathbb{R}^{n\times 1}$ the $t$-th column associated with $t$-th electrical factor. Symbol $\odot$ denotes the Hadamard product with broadcast mechanism. It is important to note that the left matrix side of the $\odot$ operation is in $\mathbb{R}^{d\times n}$, while the right matrix is in $\mathbb{R}^{n\times 1}$. The resulting matrix after the Hadamard product will also reside in the space $\mathbb{R}^{d\times n}$. At last, we use mean pooling as a readout operation to get graph-level representation.
@@ 3.3. The Implementation of Powerformer
Joint Training. To deliver the final dispatch policy based on the proposed Powerformer framework, we adopt the Dueling Deep Q Network (DQN) method [^1] The reason for selecting Dueling DQN is provide in Appendix E. Action space and reward function are placed in Appendix F. to train a deep reinforcement learning agent. Specifically, we introduce a Q-network $Q_{\theta}$ parameterized by $\theta$, which takes the state representation $\text{MFSA}(\mathcal{G})$ as input and estimates the Q-values for each possible dispatch action $d$. The dispatch policy with respect to $Q_{\theta}$ is defined as $\pi_{Q}(\mathcal{G})=\arg\max_{d\in\mathcal{D}}Q_{\theta}(\text{MFSA}(\mathcal{G}),d)$, where $\mathcal{D}$ is the dispatch action space.
@@ Appendix E The reason for selecting Dueling DQN
Appendix F Action space and reward function
Action Space is $\mathcal{A}=\{\sigma_{G_{1}}^{+},\sigma_{G_{1}}^{-},...,\sigma_{G_{N}}^{+},\sigma_{G_{N}}^{-}\}$, where $\mathcal{G}_{i}$ is the $i$-th controllable generator. Here, $\sigma_{G}^{-}$ and $\sigma_{G}^{+}$ represent the allowable decrease and increase of generation dispatch. We set $\sigma_{G}^{-}=90\%$ and $\sigma_{G}^{+}=110\%$ to follow ramp rate limits. If a generator is at its capacity limit, actions causing invalid dispatches are masked to ensure the RL agent stays within operational limits and avoids invalid configurations.
Reward Function has two main components:
1. Power Flow Component encourages maintaining the power flow within predefined transmission limits. For a single transmission interface $\Phi$, the reward penalizes deviations from the predefined limits as follows:
(18) $\mathcal{R}_{pf}(s,a;\Phi)=-|P_{\Phi}-\frac{\mathcal{P}_{\text{min},\Phi_{i}}+\mathcal{P}_{\text{max},\Phi_{i}}}{2}$
2. Economic Dispatch Component is modeled using a quadratic cost function for each generator:
(19) $\mathcal{R}_{ed}(s,a)=-\sum_{i=1}^{N_{G}}\left(\alpha_{i}(P_{i}^{G})^{2}+\beta_{i}P_{i}^{G}+\lambda_{i}\right)$
where $\alpha_{i}$, $\beta_{i}$, and $\lambda_{i}$ are cost coefficients specific to generator $i$ , and $P_{i}^{G}$ is the active power produced by that generator.
Appendix G More Ablation Studies on Three Datasets